OpenStudy (anonymous):

use the substitution to solve the following equation 4x-y-3z=-3 x-2y-2z=-3 x+y+3z=-2 a. -1 2 -1 b. 1 -2 1 c. -1 -2 1 d.1 -2 -1 NEED HEEEEELPPPPPPP

OpenStudy (anonymous):

sorry i need substitution form

OpenStudy (anonymous):

x+y+3z=-2 y=-x-3z-2

OpenStudy (anonymous):

use y=-x-3z-2 subtitution and solve

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

4x-( )-3z=-3 plug -x-3z-2 inside parenthese

OpenStudy (anonymous):

4x-y-3z=-3 4x-( -x-3z-2 )-3z=-3 4x+x+3z+2-3z=-3 5x = -3-2 x=-5/5 x=-1

OpenStudy (anonymous):

x-2(-x-3z-2)-2z=-3 x+2x+3z+4 -2z =-3 x +z =-3-4 x+z =-7

OpenStudy (anonymous):

x+y+3z=-2 -1+y+3z=-2 y + 3z = -2+1 y + 3z = -1

OpenStudy (anonymous):

-2y-2z = -2 y + 3z = -1 ( time 2) 2y+6z = -2 -2y-2z = -2 2y+6z = -2 ----------------------- Add 4z = -4 z=-1

OpenStudy (anonymous):

x+y+3z=-2 subtitution x=-1, z=-1 solve for y -1 +y+3(-1)=-2 y -3-1 = -2 y -4 = -2 y = -2+4 y = 2 the answer is a

OpenStudy (anonymous):

good night